8. A paint line can completely paint a produce with a mean time of 40 minutes with a standard deviation of 12 minutes. Nine items are selected at random to be painted.

(a) What is the probability that the mean time to paint them is (b) What is the probability that the mean time to paint the items is (c) How large a sample of items must be taken so that the sample between 35 and 45 minutes? greater than 50 minutes? mean will be within the population by less than 4 95% of the time.
Answer #1

mean = 40 and sd = 12

a)

P(35 <X < 45)

= P((35 - 40)/(12/sqrt(9)) < z < (45 - 40)/(12/sqrt(9)))

= P(-1.25 < z < 1.25)

= 0.7887

b)

P(X > 50)

= P( z > (50 - 40)/(12/sqrt(9)))

= P(z > 0.9938)

= 0.0062

c)

Given CI Level = 95%

Margin of Error(ME) = 4

std. dev. = 12

z-value of 95% CI = 1.9600

n = (z*sigma/ME)^2

n = (1.96*12/4)^2

= 35